OpenStudy (anonymous):

Please Help! Solve Systems Numerically f(x)=x^3 g(x)=4x^3-2

OpenStudy (anonymous):

What Do You Think

OpenStudy (anonymous):

I think it is gof problem but I am not sure.

OpenStudy (anonymous):

I will medal and give a testimonial

OpenStudy (anonymous):

i have no idea what this means "solve numerically" do you? are you supposed to solve \[x^3=4x^3-2\]?

OpenStudy (anonymous):

that is not so hard, you get \[3x^3=2\\ x^3=\frac{2}{3}\\ x=\sqrt[3]{\frac{2}{3}}\] but that might not be what the questions is asking

OpenStudy (anonymous):

I do not. That is also why I am stuck and confused. She said to start it off like that though so I am going to go with what you replied. Thank you so much for answering.

OpenStudy (anonymous):

I need to find y or f(x)) now. Could someone help me do that?

OpenStudy (anonymous):

since you know the value of x now just plug it back in to f(x)

OpenStudy (anonymous):

that will give you the value of f(x)

OpenStudy (anonymous):

Okay, I tried doing that and I got 2/3 so I do not know if I did that right?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

Yes! Thank you. Okay, for the remainder of the question my teacher said to check my answers by making sure if graphed it has to cross the points (2,8). What do you think she means by this..?

OpenStudy (anonymous):

cross the point (2,8)?

OpenStudy (anonymous):

Yes. She said to check my answers make sure that my graphs cross at (2,8). And if not to double check my algebra.

OpenStudy (anonymous):

She sent me this also. It is a very confusing question so I apologize, that is also why I've been stuck on it.

OpenStudy (anonymous):

well they dont cross at (2,8) is the problem

OpenStudy (anonymous):

you can use this to verify https://www.desmos.com/calculator

OpenStudy (anonymous):

and you can see in the graph she gave you they dont cross at (2,8), so im not sure where that number is coming from

OpenStudy (anonymous):

Then I don't know what she wants me to do to solve this problem really. I should be finding a point where (x,y) intersect. And thank you, I will use that. I think she was using a example to show them not crossing.

OpenStudy (anonymous):

well they cross when x = $$\sqrt[3]{\frac{2}{3}}$$ so plug that back into f(x) and g(x) and get their values, which will be the same value you already found that to be 2/3, so they cross at $$(\sqrt[3]{\frac{2}{3}}, 2/3)$$

OpenStudy (anonymous):

im not sure where the (2,8) coordinate comes from

OpenStudy (anonymous):

So then it's basically at the starting point. I will just go with those because they seem right. Me either. She just said to check your answers by seeing if my functions cross there.

OpenStudy (anonymous):

hmm ok, well when you graph it, and find the point(s) they intersect, that is solving for the system graphically when you do it with algrebra, that is solving for the system "numerically"

OpenStudy (anonymous):

if those are for sure the equations you are getting from your teacher, then that is the answer

OpenStudy (anonymous):

Yes, the for sure equations are the ones I posted first, the f(x) and g(x) ones.

OpenStudy (anonymous):

then yes your answer is the coordinate i posted

OpenStudy (anonymous):

Okay, thank you so much. :)

OpenStudy (anonymous):

you're welcome